Circular No. 7087

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

VARIABLE IN THE LARGE MAGELLANIC CLOUD A. Becker, University of Washington, on behalf of the MACHO collab-

oration (cf. IAUC 6312, plus F. Winkler and D. Paul, Middlebury College), reports the discovery of a possible nova in the LMC, located within 0′′.35 of a previously visible stellar object at α = 5^{h}35^{m}32^{s}.27, δ = −69^{o}29′52′′.1 (equinox 2000.0), which normally appears at R = 20.5, V = 21.0. This object was constant in brightness in over 900 frames taken between 1993 and 1998 at the Mt. Stromlo 1.27-m telescope. Images taken on 1998 Dec. 28.691 UT indicated a slight brightening of 0.1 mag. Sparse observations over the next 2 weeks trace a rapid increase in brightness, and the most recent observation taken on 1999 Jan. 11.742 indicates that the object is still brightening, with R = 16.9, V = 17.1. Spectra taken on Jan. 14.059 at the Cerro Tololo 1.5-m telescope by F. Winkler display narrow H emission features, plus [O III] 495.9, 500.7 nm; [N II] 654.8, 658.4 nm; and [S II] 671.7, 673.1 nm, from surrounding nebulosity evident in the images. The [S II]/[N II] ratio is anomalously high (∼ 1.5). The only spectral features that can be attributed to the brightening source are two broad absorption troughs centered at 441.8 and 513.7 nm. Finding charts can be found by

anonymous ftp to darkstar.astro.washington.edu:macho/LMC Nova.

RX J0052.1–7319 P. Kahabka, University of Amsterdam, confirms the detection of 15.3-s

pulsations for the x-ray pulsar RX J0052.1–7319 in the Small Magellanic Cloud (cf. IAUC 7081, 7082) during a 1550-s ROSAT HRI observation performed on 1996 Oct. 19.464–19.486 UT at an 8σ confidence level. The source has been detected at a large off-axis angle of 21′ with a countrate of ∼ 0.27 s^{−1} (applying a vignetting correction factor of 1.2). The observation was performed ∼ 1 month before the beginning of the 1996 Nov. 10–Dec. 9 outburst (IAUC 7081) and with a countrate that is ∼ a factor of 2 lower.

NOVA MUSCAE 1998 Visual magnitude estimates by A. Pearce, Nedlands, W. Australia:

Jan. 5.642 UT, 11.1; 6.624, 11.3; 7.62, 11.3; 8.63, 11.5; 9.60, 11.6; 10.66, 11.7; 11.79, 11.7; 12.80, 11.7; 14.79, 11.9; 16.81, 12.0.

1999 January 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green